OpenStudy (anonymous):

If ∫f(x)dx=psi(x) then ∫x^5f(x^3)dx is equal to?

OpenStudy (anonymous):

\[\int\limits_{}^{}f(x)dx=\Psi(x)\] \[\int\limits_{}^{}x^5f(x^3)dx=?\]

OpenStudy (btaylor):

I think you would integrate by parts.

OpenStudy (anonymous):

Ah how would you do that

OpenStudy (anonymous):

Oh ok i get it now

OpenStudy (anonymous):

Thanks

OpenStudy (btaylor):

you're welcome. it's a great problem.